These are systems of linear equations but it says to use linear combination and im not seeing it on the web site.
3b+2c=46
5c+b=11

Found 2 solutions by stanbon, checkley75:
Answer by stanbon(75887)

3b+2c=46
5c+b=11
--------
"linear combination" means you should solve the system by adding
or subtracting the equations.
--------------
You have:
3b+2c=46
b +5c=11
You want to eliminatie one of the variables by adding or subtracting.
You could multiply the 2nd equation by 3 to get:
3b+15c=33
Then subtract that from the 1st equation to get:
-13c = 13
c=-1
Now you can solve for b by substituting that value into one of the original equations.
b + 5(-1) = 11
b=16
==========
The system solution is c=-1 ; b=16
===========
Cheers,
Stan H.

Answer by checkley75(3666)

5C+B=11
B=-5C+11 NOW SUBSTITUTE (-5C+11) FOR B & SOLVE FOR C IN THE OTHER EQUATION.
3(-5C+11)+2C=46
-15C+33+2C=46
-13C=46-33
-13C=13
C=13/-13
C=-1 ANSWER.
5*-1+B=11
-5+B=11
B=11+5
B=16 ANSWER.
PROOF
3*16+2*-1=46
48-2=46
46=46